solve for x: 1+2sinx=0 Domain[-pi,pi]

solve for x:
1+2sinx=0
Domain[-pi,pi]
Answer #1

Salution - ceis on 1+2 Sinx =0, nefn] Implenes sinx= -1 og Sina=- sinh v Sinn = 1) or Sinx = sin( I) 0 (Because Sint-O) =-SinFor n=2, we have from @ = 27 + (-1)^(-7) = 27-I = 115 & [-7, 7] Hence 1+2 sinx=0, RE[-n, T] 57. The Completres 6 the Salution